# Anosov representations, strongly convex cocompact groups and weak eigenvalue gaps

### v3

@@ 4. The Cartan property and the uniform gap summation property
Let $\mathcal{M}$ be a compact metrizable space and $\Gamma$ a group acting on $\mathcal{M}$ by homeomorphisms. The action is called a convergence group action if for any infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ there exists a subsequence $(\gamma_{k_{n}})_{n\in\mathbb{N}}$ and $a,b\in\mathcal{M}$ such that for every compact subset $C\subset\mathcal{M}\smallsetminus\{a\}$, $\gamma_{k_{n}}|_{C}$ converges uniformly to the constant map $b$. For an infinite order element $\gamma\in\Gamma$, we denote by $\gamma^{\pm}$ the local uniform limit of the sequence $(\gamma^{\pm n})_{n\in\mathbb{N}}$. Examples of convergence group actions include the action of a non-elementary word hyperbolic group on its Gromov boundary (see [20]) and the action of a finitely generated group $\Gamma$ on its Floyd boundary $\partial_{f}\Gamma$ (see [21] and [30, Thm. 2]).
We prove a version of [12, Lem. 9.2] which shows, in many cases, that a representation $\rho$ is $P_{\theta}$-divergent when it admits a continuous $\rho$-equivariant limit map.
@@
Let $\mathcal{M}$ be a compact metrizable perfect space, $\Gamma$ a torsion free group acting on $\mathcal{M}$ by homeomorphisms and $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ a representation. Suppose that $\Gamma$ acts on $\mathcal{M}$ as a convergence group and there exists a continuous $\rho$-equivariant non-constant map $\xi:\mathcal{M}\rightarrow\mathbb{P}(\mathbb{R}^{d})$. Then for every infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ we have
$\lim_{n\rightarrow\infty}\frac{\sigma_{1}(\rho(\gamma_{n}))}{\sigma_{d-p+2}(\rho(\gamma_{n}))}=+\infty,\ \textup{where}\ p=\textup{dim}_{\mathbb{R}}\left\langle\xi\left(\mathcal{M}\right)\right\rangle.$
@@
We first prove the statement when $p=d$. If the result does not hold, then there exists $\varepsilon>0$ and a subsequence, which we continue to denote by $(\gamma_{n})_{n\in\mathbb{N}}$, such that $\frac{\sigma_{2}(\rho(\gamma_{n}))}{\sigma_{1}(\rho(\gamma_{n}))}\geqslant\varepsilon$. We may write $\rho(\gamma_{n})=k_{n}\exp(\mu(\rho(\gamma_{n}))k_{n}^{\prime}$, where $k_{n},k_{n}^{\prime}\in\mathsf{O}(d)$. Up to passing to a subsequence, there exist $\eta,\eta^{\prime}\in\mathcal{M}$ such that if $x\neq\eta^{\prime}$ then $\lim_{n}\gamma_{n}x=\eta$ and we may also assume that the sequences $(k_{n})_{n\in\mathbb{N}},(k_{n}^{\prime})_{n\in\mathbb{N}}$ converge to $k,k^{\prime}\in\mathsf{O}(d)$ respectively and $\lim_{n}\frac{\sigma_{2}(\rho(\gamma_{n}))}{\sigma_{1}(\rho(\gamma_{n}))}=C>0$. Let $x\neq\eta$ and write $\xi(x)=k_{x}P_{1}^{+}$ for some $k_{x}\in\mathsf{O}(d)$. Since $\lim_{n}\rho(\gamma_{n})\xi(x)=\xi(\eta)$, up to passing to a further subsequence, we may assume that
$\lim_{n\rightarrow\infty}\frac{\exp\big{(}\mu(\rho(\gamma_{n}))\big{)}k_{n}^{\prime}k_{x}e_{1}}{\big{|}\big{|}\exp\big{(}\mu(\rho(\gamma_{n}))\big{)}k_{n}^{\prime}k_{x}e_{1}\big{|}\big{|}}=\epsilon k^{-1}k_{\eta}e_{1}$
where $\xi(\eta)=k_{\eta}P_{1}^{+}$, $\epsilon\in\{-1,1\}$. This shows that for every $x\in\mathcal{M}$, there exits $\lambda_{x}\in\mathbb{R}$ such that:
$\big{\langle}k^{\prime}k_{x}e_{1},e_{1}\big{\rangle}=\lambda_{x}\big{\langle}k^{-1}k_{\eta}e_{1},e_{1}\big{\rangle},\ \big{\langle}k^{\prime}k_{x}e_{1},e_{2}\big{\rangle}=\lambda_{x}C^{-1}\big{\langle}k^{-1}k_{\eta}e_{1},e_{2}\big{\rangle}.$
Since $\big{\langle}\xi\left(\mathcal{M}\smallsetminus\{\eta^{\prime}\right\})\big{\rangle}=\mathbb{R}^{d}$ and $\mathcal{M}$ is perfect, there exists $x_{0}\neq\eta^{\prime}$ such that $\lambda_{x_{0}}\neq 0$. Then for every $x\neq\eta^{\prime}$ we observe that
$\big{\langle}k^{\prime}k_{x}e_{1},e_{1}\big{\rangle}=\frac{\lambda_{x}}{\lambda_{x_{0}}}\big{\langle}kk_{x_{0}}e_{1},e_{1}\big{\rangle},\ \ \big{\langle}k^{\prime}k_{x}e_{1},e_{2}\big{\rangle}=\frac{\lambda_{x}}{\lambda_{x_{0}}}\big{\langle}kk_{x_{0}}e_{1},e_{2}\big{\rangle}.$
Therefore, for every $x\neq\eta^{\prime}$, $k\xi(x)$ lies in the subspace $V=\left\langle kk_{x_{0}}e_{1}\right\rangle+e_{1}^{\perp}\cap e_{2}^{\perp}$, a contradiction since $\dim(V)\leqslant d-1$.
@@
(i) Suppose that $\rho:\Gamma\rightarrow\mathsf{SL}_{d}(\mathbb{R})$ is an irreducible representation admitting a continuous $\rho$-equivariant map $\xi:\partial_{\infty}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{d})$. Then $\rho$ is $P_{1}$-divergent and $\xi$ satisfies the Cartan property.
(ii) Suppose that $\rho^{\prime}:\Gamma\rightarrow G$ is a Zariski dense representation admitting a continuous $\rho^{\prime}$-equivariant map $\xi^{\prime}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$. Then $\rho^{\prime}$ is $P_{\theta}$-divergent and $\xi^{\prime}$ satisfies the Cartan property.
@@
Since $\rho$ is irreducible it follows by Lemma 4.5 that $\rho$ is $P_{1}$-divergent. Let $(\gamma_{n})_{n\in\mathbb{N}}$ be an infinite sequence of elements of $\Gamma$ such that $\lim_{n}\gamma_{n}=x$. By the sub-additivity of the Cartan projection $\mu$ (see [22, Fact 2.18]) and Theorem 2.5, there exists a finite subset $F$ and $C>0$ such that for every $\gamma\in\Gamma$, there exists $f\in F$ with $\big{|}\big{|}\lambda(\rho(\gamma f))-\mu(\rho(\gamma f))\big{|}\big{|}\leqslant C$. Then for large $n\in\mathbb{N}$, there exist $f_{n}\in F$ such that $\rho(\gamma_{n}f_{n})$ is $P_{1}$-proximal and
@@
(ii) Let $\tau$ be as in Proposition 2.4. Since $\rho^{\prime}$ is Zariski dense, the representation $\tau\circ\rho^{\prime}$ is irreducible. By Lemma 4.5 the representation $\tau\circ\rho^{\prime}$ is $P_{1}$-divergent and hence $\rho^{\prime}$ is $P_{\theta}$-divergent. By part (i), the $\tau\circ\rho^{\prime}$-equivariant map $\iota^{+}\circ\xi^{\prime}$ satisfies the Cartan property. It follows by Fact 4.3 that $\xi^{\prime}$ satisfies the Cartan property. ∎
We are now aiming to generalize the uniform gap summation property [22, Def. 5.2] for representations of arbitrary finitely generated groups.
@@
Let $\Gamma$ be a finitely generated group, $\rho:\Gamma\rightarrow G$ be a representation and $\theta\subset\Delta$ a finite subset of restricted roots of $G$. We say that $\rho$ satisfies the uniform gap summation property with respect to $\theta$ and the Floyd function $f:\mathbb{N}\rightarrow(0,\infty)$, if there exists $C>0$ such that
$\big{\langle}\alpha,\mu(\rho(\gamma))\big{\rangle}\geqslant-\log f(|\gamma|_{\Gamma})-C$
@@
Let $\rho:\Gamma\rightarrow G$ be a representation. If $\Gamma$ is word hyperbolic group and $\rho$ satisfies the uniform gap summation property, then it admits a pair of $\rho$-equivariant, continuous limit maps which satisfy the Cartan property (see [22, Thm. 5.3 (3)]). If $\Gamma$ is not word hyperbolic, we may similarly construct a pair of $\rho$-equivariant continuous maps from a Floyd boundary of $\Gamma$, $\partial_{f}\Gamma$, into the flag spaces $G/P_{\theta}^{+}$ and $G/P_{\theta}^{-}$. Note that when $\partial_{f}\Gamma$ is non-trivial, the action of $\Gamma$ on $\partial_{f}\Gamma$ is a convergence group action (see [30, Thm. 2]) so we obtain additional information for the action of $\rho(\Gamma)$ on its limit set in $G/P_{\theta}^{\pm}$. We prove the following lemma that we use in the following section for the proof of Theorem 1.6.
@@
Let ${\bf{p}}$ be a path in the Cayley graph of $\Gamma$ defined by the sequence of adjacent vertices $g_{0}=g,\ldots,h=g_{n}$ with $L_{f}({\bf{p}})=d_{f}(g,h)$. Since $g_{i}^{-1}g_{i+1}$ lie in a fixed generating subset of $\Gamma$, by using Proposition 4.4, we may find $C^{\prime}>0$, depending only on $\rho$, such that:
$d_{\mathbb{P}}\big{(}\Xi_{1}^{+}\big{(}\rho(g)\big{)},\Xi_{1}^{+}\big{(}\rho(h)\big{)}\big{)}$ $\leqslant\sum_{i=0}^{n-1}d_{\mathbb{P}}\big{(}\Xi_{1}^{+}\big{(}\rho(g_{i})\big{)},\Xi_{1}^{+}\big{(}\rho(g_{i+1}\big{)}\big{)}\leqslant C^{\prime}\sum_{i=0}^{n-1}\frac{\sigma_{2}(\rho(g_{i}))}{\sigma_{1}(\rho(g_{i}))}$
@@
Now define the map $\xi_{f}^{+}:\partial_{f}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{d})$ as follows: for a point $x\in\partial_{f}\Gamma$ represented by a Cauchy sequence $(\gamma_{n})_{n\in\mathbb{N}}$ with respect to the metric $d_{f}$, then
$\xi^{+}_{f}(x)=\lim_{n\rightarrow\infty}\Xi_{1}^{+}\big{(}\rho(\gamma_{n})\big{)}$
The previous inequality shows that the limit $\lim_{n}\Xi_{1}^{+}\big{(}\rho(\gamma_{n})\big{)}$ is independent of the choice of the sequence $(\gamma_{n})_{n\in\mathbb{N}}$ representing $x$, since for any other sequence $(\gamma_{n}^{\prime})_{n\in\mathbb{N}}$ with $x=\lim_{n}\gamma_{n}^{\prime}$, we have $\lim_{n}d_{f}(\gamma_{n},\gamma_{n}^{\prime})=0$. Finally, $\xi^{+}_{f}$ is well defined and Lipshitz by (4), and hence continuous. By identifying $G/P_{\theta}^{-}$ with $G/P_{\theta^{\star}}^{+}$, we similarly obtain the limit map $\xi_{f}^{-}$.
Suppose that $\rho(\Gamma)$ is $P_{1}$-proximal. By the definition of the map $\xi_{f}^{+}$, if $\rho(\gamma_{0})$ is $P_{1}$-proximal, then $\xi_{f}^{+}(\gamma_{0}^{+})$ is the unique attracting fixed point of $\rho(\gamma_{0})$ in $\mathbb{P}(\mathbb{R}^{d})$. Since $\Gamma$ acts minimally on $\partial_{f}\Gamma$ we have $\xi_{f}^{+}(\partial_{f}\Gamma)=\Lambda_{\Gamma}$. ∎

### v4

@@ 4. The Cartan property and the uniform gap summation property
Let $\mathcal{M}$ be a compact metrizable space and $\Gamma$ a group acting on $\mathcal{M}$ by homeomorphisms. The action is called a convergence group action if for any infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ there exists a subsequence $(\gamma_{k_{n}})_{n\in\mathbb{N}}$ and $x,y\in\mathcal{M}$ such that for every compact subset $\mathcal{C}\subset\mathcal{M}\smallsetminus\{x\}$, $\gamma_{k_{n}}|_{\mathcal{C}}$ converges uniformly to the constant map $y$. For an infinite order element $\gamma\in\Gamma$, we denote by $\gamma^{\pm}$ the local uniform limit of the sequence $(\gamma^{\pm n})_{n\in\mathbb{N}}$. Examples of convergence group actions include the action of a non-elementary word hyperbolic group on its Gromov boundary (see [22]) and the action of a finitely generated group $\Gamma$ on its Floyd boundary $\partial_{f}\Gamma$ (see [23] and [32, Thm. 2]).
We prove a version of [14, Lem. 9.2] which shows, in many cases, that a representation $\rho$ admitting a continuous $\rho$-equivariant limit map in the flag space $G/P_{\theta}^{+}$ is $\theta$-divergent. For a subset $\mathcal{C}\subset\mathbb{P}(\mathbb{R}^{d})$ define $\langle\mathcal{C}\rangle:=\textup{span}\big\{v\in\mathbb{R}^{d}\smallsetminus\{{\bf 0}\}:[v]\in\mathcal{C}\big\}$. We shall prove first the following lemma.
@@
Let $\mathcal{M}$ be a compact metrizable perfect space, $\Gamma$ a torsion-free group acting on $\mathcal{M}$ by homeomorphisms and $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ a representation. Suppose that $\Gamma$ acts on $\mathcal{M}$ as a convergence group and there exists a continuous $\rho$-equivariant non-constant map $\xi:\mathcal{M}\rightarrow\mathbb{P}(\mathbb{R}^{d})$. Then for every infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ we have
$\lim_{n\rightarrow\infty}\frac{\sigma_{1}(\rho(\gamma_{n}))}{\sigma_{d-p+2}(\rho(\gamma_{n}))}=+\infty$
where $p=\textup{dim}_{\mathbb{R}}\left\langle\xi\left(\mathcal{M}\right)\right\rangle$.
@@
We first prove the statement when $p=d$.
If the result does not hold, then there exists $\varepsilon>0$ and a subsequence, which we continue to denote by $(\gamma_{n})_{n\in\mathbb{N}}$, such that $\frac{\sigma_{2}(\rho(\gamma_{n}))}{\sigma_{1}(\rho(\gamma_{n}))}\geqslant\varepsilon$. We may write $\rho(\gamma_{n})=k_{n}\exp(\mu(\rho(\gamma_{n}))k_{n}^{\prime}$, where $k_{n},k_{n}^{\prime}\in\mathsf{O}(d)$. Up to passing to a subsequence, there exist $\eta,\eta^{\prime}\in\mathcal{M}$ such that if $x\neq\eta^{\prime}$ then $\lim_{n}\gamma_{n}x=\eta$ and we may also assume that the sequences $(k_{n})_{n\in\mathbb{N}},(k_{n}^{\prime})_{n\in\mathbb{N}}$ converge to $k,k^{\prime}\in\mathsf{O}(d)$ respectively, $\lim_{n}\frac{\sigma_{2}(\rho(\gamma_{n}))}{\sigma_{1}(\rho(\gamma_{n}))}=C>0$ and for every $i>1$ the limit $\lim_{n}\frac{\sigma_{i}(\rho(\gamma_{n}))}{\sigma_{1}(\rho(\gamma_{n}))}$ exists.
For $z\in\partial_{\infty}\Gamma$ write $\xi(x)=k_{z}P_{1}^{+}$ for some $k_{z}\in\mathsf{O}(d)$. Now let $x\in\partial_{\infty}\Gamma\smallsetminus\{\eta\}$. Since $\lim_{n}\rho(\gamma_{n})\xi(x)=\xi(\eta)$, up to passing to a further subsequence, we may assume that
[…]
$\big\langle k^{\prime}k_{x}e_{1},e_{1}\big\rangle$ $=\lambda_{x}\big\langle k^{-1}k_{\eta}e_{1},e_{1}\big\rangle$
$\big\langle k^{\prime}k_{x}e_{1},e_{2}\big\rangle$ $=\lambda_{x}C^{-1}\big\langle k^{-1}k_{\eta}e_{1},e_{2}\big\rangle.$
Since $\mathcal{M}$ is perfect we have $\big\langle\xi\left(\mathcal{M}\smallsetminus\{\eta^{\prime}\right\})\big\rangle=\mathbb{R}^{d}$ and also there exists $x_{0}\neq\eta^{\prime}$ with $\lambda_{x_{0}}\neq 0$. Then for every $x\neq\eta^{\prime}$ we observe that
$\big\langle k^{\prime}k_{x}e_{1},e_{1}\big\rangle$ $=\frac{\lambda_{x}}{\lambda_{x_{0}}}\big\langle kk_{x_{0}}e_{1},e_{1}\big\rangle$
$\big\langle k^{\prime}k_{x}e_{1},e_{2}\big\rangle$ $=\frac{\lambda_{x}}{\lambda_{x_{0}}}\big\langle kk_{x_{0}}e_{1},e_{2}\big\rangle.$
Therefore, for every $x\neq\eta^{\prime}$, $k\xi(x)$ lies in the subspace $V=\left\langle kk_{x_{0}}e_{1}\right\rangle+e_{1}^{\perp}\cap e_{2}^{\perp}$, a contradiction since $\dim(V)\leqslant d-1$. This completets the proof when $p=d$.
@@
  1. (i)
    Suppose that $\rho:\Gamma\rightarrow\mathsf{SL}_{d}(\mathbb{R})$ is an irreducible representation admitting a continuous $\rho$-equivariant map $\xi:\partial_{\infty}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{d})$. Then $\rho$ is $P_{1}$-divergent and $\xi$ satisfies the Cartan property.
  2. (ii)
    Suppose that $\rho^{\prime}:\Gamma\rightarrow G$ is a Zariski dense representation admitting a continuous $\rho^{\prime}$-equivariant map $\xi^{\prime}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$. Then $\rho^{\prime}$ is $P_{\theta}$-divergent and $\xi^{\prime}$ satisfies the Cartan property.
@@
[…]
@@
(ii) Let $\tau_{\theta}:G\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ and $(\iota^{+},\iota^{-})$ be as in Proposition 2.4. Since $\rho^{\prime}$ is Zariski dense, the representation $\tau_{\theta}\circ\rho^{\prime}$ is irreducible. By Lemma 4.5 the representation $\tau_{\theta}\circ\rho^{\prime}$ is $P_{1}$-divergent and hence $\rho^{\prime}$ is $P_{\theta}$-divergent. By part (i), the $\tau_{\theta}\circ\rho^{\prime}$-equivariant map $\iota^{+}\circ\xi^{\prime}$ satisfies the Cartan property. It follows by Fact 4.3 that $\xi^{\prime}$ satisfies the Cartan property. ∎
We are now aiming to generalize the uniform gap summation property [24, Def. 5.2] for representations of arbitrary finitely generated groups.
@@
Let $\Gamma$ be a finitely generated group, $\rho:\Gamma\rightarrow G$ a representation and $\theta\subset\Delta$ a finite subset of restricted roots of $G$. We say that $\rho$ satisfies the uniform gap summation property with respect to $\theta$ and the Floyd function $f:\mathbb{N}\rightarrow(0,+\infty)$, if there exists $C>0$ such that
$\alpha\big(\mu(\rho(\gamma))\big)\geqslant-\log f(|\gamma|_{\Gamma})-C$
@@
Let $\rho:\Gamma\rightarrow G$ be a representation. If $\Gamma$ is word hyperbolic group and $\rho$ satisfies the uniform gap summation property, then it admits a pair of $\rho$-equivariant, continuous limit maps which satisfy the Cartan property (see [24, Thm. 5.3 (3)]). If $\Gamma$ is not word hyperbolic, we may similarly construct a pair of $\rho$-equivariant continuous maps from a Floyd boundary of $\Gamma$, $\partial_{f}\Gamma$, into the flag spaces $G/P_{\theta}^{+}$ and $G/P_{\theta}^{-}$. Note that when $\partial_{f}\Gamma$ is non-trivial, the action of $\Gamma$ on $\partial_{f}\Gamma$ is a convergence group action (see [32, Thm. 2]) so we obtain additional information for the action of $\rho(\Gamma)$ on its limit set in $G/P_{\theta}^{+}$. For a subgroup $H$ of $G$ containing a $P_{\theta}$-proximal element, its limit set in $G/P_{\theta}^{+}$ is the closure of attracting fixed points of $\theta$-proximal elements of $H$. In the case where $H\subset\mathsf{SL}_{d}(\mathbb{R})$ and $G/P_{\theta}^{+}=\mathbb{P}(\mathbb{R}^{d})$, $P_{\theta}^{+}=P_{1}$, we denote by $\Lambda_{H}$ its proximal limit set in $\mathbb{P}(\mathbb{R}^{d})$. If $H$ is in addition an irreducible subgroup of $\mathsf{GL}_{d}(\mathbb{R})$, then $H$ acts minimally on $\Lambda_{H}$, see [6, Lem. 2.5].
We prove the following lemma that we use in the following section for the proof of Theorem 1.6.
@@
Let ${\bf{p}}\subset C_{\Gamma}$ be a path in the Cayley graph of $\Gamma$ defined by the sequence of adjacent vertices $g_{0}=g,\ldots,h=g_{n}$ with $L_{f}({\bf{p}})=d_{f}(g,h)$. Since for every $i$, $g_{i}^{-1}g_{i+1}$ lies in a fixed generating subset of $\Gamma$, by Proposition 4.4, there is $C^{\prime}>0$, depending only on $\rho$, such that:
$d_{\mathbb{P}}\big(\Xi_{1}^{\pm}\big(\rho(g)\big),\Xi_{1}^{\pm}\big(\rho(h)\big)\big)$ $\leqslant\sum_{i=0}^{n-1}d_{\mathbb{P}}\big(\Xi_{1}^{\pm}\big(\rho(g_{i})\big),\Xi_{1}^{\pm}\big(\rho(g_{i+1}\big)\big)\leqslant C^{\prime}\sum_{i=0}^{n-1}\frac{\sigma_{2}(\rho(g_{i})^{\pm 1})}{\sigma_{1}(\rho(g_{i})^{\pm 1})}$
@@
Now define the maps $\xi_{f}^{+}:\partial_{f}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{d})$ and $\xi_{f}^{-}:\partial_{f}\Gamma\rightarrow\mathsf{Gr}_{d-1}(\mathbb{R}^{d})$ as follows: for a point $x\in\partial_{f}\Gamma$ represented by a Cauchy sequence $(\gamma_{n})_{n\in\mathbb{N}}$ with respect to the metric $d_{f}$, define $\xi_{f}^{\pm}(x)$,
$\xi^{\pm}_{f}(x):=\lim_{n\rightarrow\infty}\Xi_{1}^{\pm}\big(\rho(\gamma_{n})\big)$
The bound (4) shows that the limit $\lim_{n}\Xi_{1}^{+}\big(\rho(\gamma_{n})\big)$ is independent of the choice of the sequence $(\gamma_{n})_{n\in\mathbb{N}}$ representing $x$, since for any other sequence $(\gamma_{n}^{\prime})_{n\in\mathbb{N}}$ with $x=\lim_{n}\gamma_{n}^{\prime}$, we have $\lim_{n}d_{f}(\gamma_{n},\gamma_{n}^{\prime})=0$. Finally, $\xi^{+}_{f}$ is well-defined and Lipschitz by (4) and hence continuous. By identifying $G/P_{\theta}^{-}$ with $G/P_{\theta^{\star}}^{+}$, we similarly obtain deduce that the limit map $\xi_{f}^{-}$ is well-defined and continuous.
Suppose that $\rho(\Gamma)$ is $P_{1}$-proximal. By the definition of the map $\xi_{f}^{+}$ (resp. $\xi_{f}^{-}$), if $\rho(\gamma_{0})$ is $P_{1}$-proximal (resp. $P_{d-1}$-proximal), then $\xi_{f}^{+}(\gamma_{0}^{+})$ (resp. $\xi_{f}^{-}(\gamma_{0}^{+})$) is the unique attracting fixed point of $\rho(\gamma_{0})$ in $\mathbb{P}(\mathbb{R}^{d})$ (resp. $\mathsf{Gr}_{d-1}(\mathbb{R}^{d})$). Since $\Gamma$ acts minimally (e.g. see [32]) on $\partial_{f}\Gamma$ we deduce that $\xi_{f}^{+}(\partial_{f}\Gamma)$ is the proximal limit set of $\Gamma$ in $\mathbb{P}(\mathbb{R}^{d})$. ∎